Find the present value of the future amount. Assume 365 days in a year. Round to the nearest cent.
$16,000 for 10 months; money earns 9%

Answer by rfer(16322)

A=16000(1+0.09/(12)(10))^(10/12)
A=16000*(1.075)^0.8333
A=16000*1.062
A=$16,992.00